Hi. I have taken a certification test this past weekend and want to know if I got the answer to the problem below correct or not. Thanks!
(3+1)
______
(6+4)^2

Answer by checkley71(8403)

(3+1)/(6+4)^2=4/10^2=4/100=1/25 OR .04 ANSWER.